OpenStudy (anonymous):

Simplify the given expression. (−3 + 2i) • (2 + i) −8 + i −6 + 3i −3 + 4i −1 + 2i

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (mathstudent55):

Use FOIL to multiply the two numbers together. Remember that \(i^2 = -1\).

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

your answer is going to be -6+3i

OpenStudy (mathstudent55):

(−3 + 2i) • (2 + i) Use FOIL: = (-3)(2) + (-3)(i) + (2i)(2) + (2i)(i) = -6 - 3i + 4i + 2i^2 = -6 + i + 2(-1) = -6 + i - 2 = -8 + i

OpenStudy (anonymous):

@HelpBlahBlahBlah no Im not named after the princess aurora My mother named me that because We are mexican and she thought that it was a pretty name for a mexcian girl

OpenStudy (anonymous):

@HelpBlahBlahBlah im sorry are you a girl or boy

OpenStudy (helpblahblahblah):

why?

OpenStudy (anonymous):

I dont know that was just a question that i ask

OpenStudy (helpblahblahblah):

oh okay

OpenStudy (mathstudent55):

I think it's funny that @aurora007 already got two medals for her answer. Has anyone noticed it's wrong?

OpenStudy (anonymous):

@HelpBlahBlahBlah where are you from

OpenStudy (helpblahblahblah):

from the world of course

OpenStudy (anonymous):

i know that much but i was wondering like what state or country are you from

OpenStudy (helpblahblahblah):

from some where

OpenStudy (helpblahblahblah):

okay please leave me alone

OpenStudy (anonymous):

mathstudent55 are you the one that said something stupid about Aurora007

OpenStudy (mathstudent55):

I did write something about aurora's answer to the math question. I didn't write anything stupid about aurora.

OpenStudy (anonymous):

next time dont write anything at all okay and your just( muy loco)

OpenStudy (mathstudent55):

Ok, do you think it's better for the person who asked the question to have a wrong answer?

OpenStudy (anonymous):

I just got home and my cousin was on my laptop and I didnt even answer the question she did and my abuela was suppose to be watching her but she still ended up with my laptop anyway

OpenStudy (mathstudent55):

People ask math questions here to get help with their homework. How does a wrong answer help? It just confuses more the person who needs help.

OpenStudy (anonymous):

i am a working person i told you already i just got home and i didnt anwser the question okay

OpenStudy (mathstudent55):

If what you're saying is true, then you shouldn't let other people use your account and make you look bad.

OpenStudy (mathstudent55):

Still, that's not the point. The point I was making is that you received two medals for giving out an incorrect answer. I think that is funny.

OpenStudy (anonymous):

If I was at work how would i know if someone has been in my room on my stuff i wouldnt because i was gone all freaking day and your really pissing me off because i told you already that my cousin had my damn laptop and she answerd the question okay not me

OpenStudy (mathstudent55):

I find it interesting that one moment your prima is using your computer at your casa and writes an incorrect answer, and just 4 minutes later, you are using the computer and explaining how your name came about. You sure drive fast from el trabajo to la casa. By the way, I think Aurora is a lovely name. It means Dawn.

OpenStudy (anonymous):

it is also interesting how her first question was an hour ago, and yet she says she was gone all day.

OpenStudy (mathstudent55):

Ok, Aurora, I'm not picking on you. We were just having some fun. Sorry if you got upset. I'll stop now. Have a great day.

OpenStudy (anonymous):

ellos listen I work 30 min away from my casa and my prima was on my computadora so she decided to use my laptop for her homework and she needed help so she found this website that i used to use when i was still in la escuela my aubela was watching her while i was at work

OpenStudy (anonymous):

mathstudent55 no your okay im just letting you know what happened

OpenStudy (anonymous):

abracadabra you are a eldesgraciado to me

OpenStudy (helpblahblahblah):

thats rude

OpenStudy (anonymous):

you know what i said

OpenStudy (helpblahblahblah):

i speak spanish and what you said was not very nice >_< especially to my friend

OpenStudy (anonymous):

bien dile a tu amigo que no me gusta la forma en que piensas que te di una respuesta a su pregunta

OpenStudy (anonymous):

you can tell me yourself @aurora007

OpenStudy (helpblahblahblah):

leave me alone tell her by your self i dont like rude people especcially when they do tuff to my friend

OpenStudy (helpblahblahblah):

and apply it to your self to

OpenStudy (anonymous):

ustedes son tan loco i dont care para ustedes realmente he usado este sitio cuando yo estaba en la escuela no im en la escuela ya soy un estudiante universitario que acaba de conseguir de trabajo y yo no aprecio lo que ustedes están hablando a mí como soy loco o un cierto tipo de persona estúpida

OpenStudy (helpblahblahblah):

no estoy hablando asi eres tu

OpenStudy (anonymous):

@HelpBlahBlahBlah is not crazy, you were mean to her first.

OpenStudy (helpblahblahblah):

thanks @abracadabra

OpenStudy (mathstudent55):

I think we should all be nice to each other. Let's get back to math, and stop using the post area to call each other names. Aurora, it was nice meeting you online. If you ever need help with math, and I know the material, I'll try to help. You can tag me any time. To @HelpBlahBlahBlah and @abracadabra , the same applies. Let's talk to each other nicely using messages, not the posts. If you have questions on math, tag me. I'll do my best to help.

OpenStudy (helpblahblahblah):

okay @mathstudent55

OpenStudy (anonymous):

alright @mathstudent55

OpenStudy (helpblahblahblah):

it was nice metting you to @aurora007

OpenStudy (mathstudent55):

And to think that this whole thing started with a simple question on multiplying complex numbers...

OpenStudy (anonymous):

tengo que ir a buscar a la cama tengo que levantarme a las tres de la mañana para el trabajo

OpenStudy (mathstudent55):

Good night everyone. gtg bye

OpenStudy (helpblahblahblah):

goodnight

OpenStudy (anonymous):

good night @mathstudent55

OpenStudy (anonymous):

HelpBlahBlahBlah ,abracadabra ,@mathstudent55 buenos noches

OpenStudy (helpblahblahblah):

buenas noches

OpenStudy (anonymous):

Encantada de conocerte también , abracadabra ( nice to meet you too)

OpenStudy (helpblahblahblah):

perdone por la molestia

OpenStudy (anonymous):

esta bien yo sólo sé la próxima vez para encerrarme en mi habitación así que mi primo no puede entrar nunca más

OpenStudy (helpblahblahblah):

okay dios te bendiga

OpenStudy (anonymous):

dios bendiga a usted también tiene un buen día

OpenStudy (helpblahblahblah):

lo mismo

OpenStudy (anonymous):

yo no creo que su amigo quería de hablar conmigo tan le dicen que dije que pido disculpas por toda la situación y los nombres ofensivos que la llamé y le diga que dije fue un placer conocerte ella y que cosas no se deja llevar un poco y voy a pedir disculpas a mí ya mi nombre primo

OpenStudy (helpblahblahblah):

es queella esta comiendo por eso no te contesta

OpenStudy (anonymous):

gracias por lo mucho

OpenStudy (helpblahblahblah):

de nada

OpenStudy (anonymous):

usted es una persona muy agradable

OpenStudy (helpblahblahblah):

gracias

OpenStudy (anonymous):

de nada Lo siento si te he ofendido o su amigo hoy

OpenStudy (helpblahblahblah):

its okay no offendistes i got a little mad tho

OpenStudy (anonymous):

ENTIENDO

OpenStudy (helpblahblahblah):

okay adios